# Multi-level protein pre-training with Vabs-Net

### v3

@@ 1 Introduction
We designe a series of downstream tasks to assess the impact of Vabs-Net. These tasks include Enzyme Commission (EC) number prediction and Gene Ontology (GO) term prediction, which focus on the global properties of proteins, as well as binding site prediction, which emphasizes protein local properties. Furthermore, we incorporated molecular docking, a task known to rely on atom modeling of proteins, as each atom may interact with small molecules. We incorporate the atom-level representation generated by our model and input it into an existing docking model, subsequently evaluating its effectiveness through performance improvement in molecular docking tasks. Across these tasks, our model, Vabs-Net, significantly outperforms previous baselines, demonstrating its superior efficacy.
In summary, our contributions are as follows:
@@
Figure 1: With all-atom added, the possible range for residue position is limited, thus resulting in easier prediction for residue position and angle between edges, etc. When all atoms are utilized, the prediction of residue positions and inter-edge angles relies predominantly on other atoms rather than residues themselves. This may result in a suboptimal representation of residues.
@@ 2 Related Works
Pre-trained protein models. Pre-training using protein sequences has recently attracted significant attention due to its potential applications in areas such as protein structure prediction and drug discovery (Lin et al., 2023; Chen et al., 2023a). A family of models, which includes ProtTrans (Elnaggar et al., 2020), ESM-1b (Rao et al., 2021), and ESM2 (Lin et al., 2023), utilize individual protein sequences as input and undergo pre-training by optimizing the masked language model objective based on encoder-only architectures. A prompt-based pre-trained protein model (Wang et al., 2022b) has been introduced to guide multi-task pre-training. Additionally, xTrimoPGLM (Chen et al., 2023a) is proposed to further explore the potential of a unified pre-training strategy that could comprehensively characterize the underlying distributions of protein sequences.
Recent advancements in structure-based pre-training models have demonstrated significant improvements in performance across a diverse range of downstream tasks. The prevailing protein pre-training models primarily focus on predicting a diverse range of physical quantities, including torsion angles, angles between edges, alpha carbon positions, and distances, as well as residue types. (Zhang et al., 2022b; Guo et al., 2022; Chen et al., 2023b; Wang et al., 2022b). Contrastive learning has demonstrated exceptional performance in a recent study based on GearNet (Zhang et al., 2022b). Additionally, recent attempts for pre-training with protein surfaces have been made (Lee et al., 2023). Numerous diffusion-based models have been proposed (Huang et al., 2023; Zhang et al., 2023). For instance, SiamDiff (Zhang et al., 2023) employs a pre-trained protein encoder through sequence-structure joint diffusion modeling.
@@ 3.1 Vector Aware Bilevel Sparse Attention Network
Protein graph construction. We first construct both residue-level and atom-level graphs. A residue level graph can be defined as $\mathcal{G}_{res}=(\mathcal{V}_{CA},\mathcal{E}_{res})$, where $\mathcal{V}_{CA}$ includes all alpha carbon nodes and $\mathcal{E}_{res}$ includes all residue edges. An atom level graph can be defined as $\mathcal{G}_{atom}=(\mathcal{V},\mathcal{E}_{atom})$, where $\mathcal{V}$ includes all-atom nodes, i.e. $\mathcal{V}_{CA}\subset\mathcal{V}$ and $\mathcal{E}_{atom}$ includes all atom level edges. The construction of edges involves the selection of k nearest neighbors in the spatial context for both residue and atom-level graph nodes. The residue-level graph shares alpha carbon nodes with the atom-level graph. Consequently, the bilevel graph of protein can be written as $\mathcal{G}=(\mathcal{V},\mathcal{E}_{res},\mathcal{E}_{atom})$. The model architecture can be seen in Figure2. This approach facilitates information exchange between the atom and residue levels, thereby not only allowing atoms to acquire an expanded receptive field, but also enabling residues to obtain detailed structural information.
To comprehensively encapsulate global information on protein, we introduce a virtual origin point positioned at geometry center. This point is connected with every atom and residue, integrating overall protein representation.
Node Encoding. In our model, node embeddings incorporate both atom type and residue type. Additionally, we utilize the Large language model (LLM) version ESM (Lin et al., 2023) to leverage sequence information. The representation of atom nodes can be seen as follows:
@@
where $ATOM$ and $RES$ stand for the type of atom and residue. $W_{E}$ maps the dimension of ESM feature to node dimension.
Edge Distance Encoding. To encode the distance between atoms or residues, we use the Gaussian kernel.
@@
Edge Vector Encoding. In previous protein-pretraining modeling, structural information is traditionally encoded using the distance between residues. While this distance-based structural encoding may suffice for residue modeling, encoding distance is not informative enough for atom modeling since the number of atoms is much larger than that of the residue. We find that encoding edge directions in both the residue local coordinate system and the absolute global coordinate system can alleviate this problem. The residue local coordinate system can be formulated as following steps with three backbone atoms(CA, C, N) in each residue. We use $\boldsymbol{r}$ to indicate the position of atoms and set $\boldsymbol{v}_{N}=\boldsymbol{r}_{N}-\boldsymbol{r}_{CA}$ and $\boldsymbol{v}_{C}=\boldsymbol{r}_{C}-\boldsymbol{r}_{CA}$ as the vector of edges between N and CA, CA and C in the absolute global coordinate system. More generally, We define $\boldsymbol{v}_{l}$ and $\boldsymbol{v}_{g}=\boldsymbol{r}_{atom}-\boldsymbol{r}_{CA}$ are the vectors of edges from CA to an atom in the same residue in the local residue coordinate system and global coordinate system, respectively. They serve to encode edge directions in subsequent steps. Then we can get a local coordinate system with a rotational matrix $\boldsymbol{R}$, which is used to transform vectors in global coordinates to local coordinates. $\boldsymbol{R}$ is constructed with unit vector $\boldsymbol{u},\boldsymbol{v},\boldsymbol{w}$ in XYZ axis of residue local coordinate.
@@
where $\boldsymbol{l}(i-j)$ is the learnable positional encoding based on the sequence distance between two residues. $\boldsymbol{W_{g}}$ and $\boldsymbol{W_{f}}$ maps dimension of distance and direction encoding to dimension of edge.
Sparse Attention Module. Numerous pre-training models for molecules based on Graphormer have achieved state-of-the-art performance in downstream tasks (Zhou et al., 2023). However, a fully connected Graphormer model can pose significant computational challenges (Lu et al., 2023). To tailor Graphormer’s attention mechanism to the all-atom level graph, we introduce a Sparse Graph Attention Module aimed at reducing computational load and preventing overfitting in downstream tasks. The sparse attention module is formulated as follows:
@@
where $\boldsymbol{x}^{l}_{i}$ is the representation of node $i$ in the $l$th layer. $a^{l}_{i,j}$ is the attention weight. Set $\mathcal{N}_{i}$ includes all nodes connected to node $i$. $\boldsymbol{W_{B}}$ is a linear layer for bias calculation. $\boldsymbol{e}_{i,j}$ is the edge representation between node $i$ and node $j$. For simplicity, the $h$ dimension in multi-head attention is omitted.
@@
Span mask protein chain(SMPC) residue type prediction. During the pre-training phase, span masking is applied to mask residue types of consecutive biologically meaningful substructures. Specifically, only the alpha carbon atom is retained for the span-masked residues, while all other atoms are eliminated. This aims to prevent information leakage from side chain atoms and other backbone atoms, which could make residue type prediction and torsion angle prediction trivial. We find even backbone atoms can leak residue type to some extent. Subsequently, only alpha carbon representation is employed for residue type prediction. Cross-entropy loss serves as the objective function for predicting residue types.
Span mask protein chain(SMPC) side chain and backbone torsion angle prediction. Drawing upon the feature of alpha carbon atoms in SMPC part, Vabs-Net predicts the cosine and sine components for seven side chain and backbone torsion angle angles. In addition to the L1 norm loss for cosine and sine values, the loss function encompasses a term that ensures the normalization of the sum of squared sine and cosine values to promote accurate angular representation. The loss function for this task is similar to the one used in AlphaFold2 (Jumper et al., 2021).
Atom position and distance prediction. Apart from the residue level pre-training tasks. To learn atom-level structural information, We randomly choose consecutively residues independent of SMPS. For these residues, we simply add Gaussian noise to atoms in those residues instead of removing atoms. For these atoms, positions are predicted by a movement prediction head.
@@
Solvent Accessible Surface Area(SASA) prediction. SASA describes the surface area of a biomolecule, such as a protein or nucleic acid, that is accessible to a solvent. It could be used to understand the shape of a protein. We used freesasa (Mitternacht, 2016) to calculate sasa of each atom. The loss function for SASA prediction is L1 loss.
@@ 4 Experiments
In this section, we report our experiment setup and results in training and evaluation of our models for pre-training and downstream tasks. Detailed information about downstream task settings can be seen in Appendix C.
@@
Method
BP
MF
CC
EC
DEEPFRI(Gligorijević et al., 2021)
0.399
0.465
0.460
0.631
MTL(Wang et al., 2022b)
0.445
0.640
0.503
0.869
GRADNORM(Chen et al., 2018; Wang et al., 2022b)
0.466
0.643
0.504
0.874
LM-GVP(Wang et al., 2022a)
0.417
0.545
0.527
0.664
ROTOGRAD(Javaloy & Valera, 2022)
0.470
0.638
0.509
0.876
CDConv(Fan et al., 2022)
0.453
0.654
0.479
0.820
ESM_1b(Lin et al., 2023)
0.470
0.657
0.488
0.864
GearNet(Zhang et al., 2022b)
0.490
0.650
0.486
0.874
Siamdiff(Zhang et al., 2023)
0.857
GearNet-ESM(Zhang et al., 2022b)
0.516
0.684
0.506
0.890
SiamDiff-ESM(Zhang et al., 2023)
0.897
GearNet-ESM-INR-MC(Lee et al., 2023) 0.518
0.683
0.504
0.896
Vabs-Net-no-SMPC
0.512
0.697 0.573 0.898
Vabs-Net 0.542 0.708 0.610 0.914
@@ 4.3 Protein binding site prediction
Method pre-train
B277
DT198
ASTEX85
CHEN251
COACH420
FPocket(Le Guilloux et al., 2009) $\times$
31.5
23.2
34.1
25.4
30.0
SiteHound(Hernandez et al., 2009) $\times$
36.4
23.1
38.9
29.4
34.9
MetaPocket2(Macari et al., 2019) $\times$
37.3
25.8
37.5
32.8
37.7
DeepSite(Jiménez et al., 2017) $\times$
34.0
29.1
37.4
27.4
33.9
P2Rank(Krivák & Hoksza, 2018) $\times$
49.8
38.6
47.4
56.5
45.3
ESM2_150M(Lin et al., 2023) $\surd$
19.6
16.6
20.5
18.9
22.0
GearNet(Zhang et al., 2022b) $\surd$
39.9
35.8
41.0
36.4
41.3
Siamdiff(Zhang et al., 2023) $\surd$
37.7
31.0
40.7
35.3
40.3
Vabs-Net $\times$
57.7
48.6
57.8
53.2
61.4
Vabs-Net $\surd$ 60.1 52.0 58.8
56.3
64.1
@@
Method pre-train
auc
TargetDNA(Hu et al., 2016) $\times$
0.825
DNAPred(Zhu et al., 2019) $\times$
0.845
SVMnuc(Su et al., 2019) $\times$
0.812
COACH-D(Wu et al., 2018) $\times$
0.761
NucBind(Su et al., 2019) $\times$
0.797
DNABind(Liu & Hu, 2013) $\times$
0.858
@@
ESM2_150M(Lin et al., 2023) $\surd$
0.779
GearNet(Zhang et al., 2022b) $\surd$
0.849
Siamdiff(Zhang et al., 2023) $\surd$
0.823
Vabs-Net $\times$
0.912
@@
Setting We have construct a large high-quality small molecule binding site dataset, and we use this dataset to train our model and other baselines, which will be open access to the public. For all-atom models, the label of Alpha C is the residue label. In DNA and RNA binding site prediction, we use the AUC for evaluation. Baselines are finetuned using the same setting as their original paper.
In the preparation of our small molecule training dataset, we utilize three distinct datasets: CASF-2016 coreset, PDBBind v2020 refined set, and MOAD. During dataset preparation, we follow the original ligand records in these datasets to extract the protein and ligand components based on the corresponding PDB ID from RCSB. The extracted segments undergo a structural repair process using in-house scripts. For proteins, this process includes the repair of missing residues, replenishment of absent heavy atoms, and addition of hydrogen atoms. In the case of ligands, the process involves repairing bond orders, adding hydrogen atoms, and determining the correct protonation states based on the pocket environment. A comparison of our dataset and frequently used scPDB can be seen in Appendix A. To avoid leakage of information, MMSeqs2 is used to filter out high protein sequence similarity(similarity above 40% which is also used in AlphaFold2 (Jumper et al., 2021)) with test sets.

### v4

# Pre-Training Protein Bi-level Representation Through Span Mask Strategy On 3D Protein Chains

@@ 1 Introduction
We design a series of downstream tasks to assess the impact of Vabs-Net. These tasks include Enzyme Commission (EC) number prediction and Gene Ontology (GO) term prediction, which focus on the global properties of proteins, as well as binding site prediction, which emphasizes protein local properties. Furthermore, we incorporated molecular docking, a task known to rely on atom modeling of proteins, as each atom may interact with small molecules. We incorporate the atom-level representation generated by our model and input it into an existing docking model, subsequently evaluating its effectiveness through performance improvement in molecular docking tasks. Across these tasks, our model, Vabs-Net, significantly outperforms previous baselines, demonstrating its superior efficacy. In summary, our contributions are as follows:
@@
Figure 1: With all-atom added, the possible range for residue position is limited, thus resulting in easier prediction for residue position and angle between edges, etc. When all atoms are utilized, the prediction of residue positions and inter-edge angles relies predominantly on other atoms rather than residues themselves.
@@ 2 Related Works
Pre-trained Protein Models Pre-training using protein sequences has recently attracted significant attention due to its potential applications. A family of models, which includes ProtTrans (Elnaggar et al., 2020), ESM-1b (Rao et al., 2021), and ESM2 (Lin et al., 2023), utilize individual protein sequences as input and undergo pre-training by optimizing the masked language model objective. A prompt-based pre-trained protein model (Wang et al., 2022b) has been introduced to guide multi-task pre-training. Additionally, xTrimoPGLM (Chen et al., 2023a) is proposed to further explore the potential of a unified pre-training strategy.
Recent advancements in structure-based pre-training models have demonstrated significant improvements in performance across a diverse range of downstream tasks. The prevailing protein pre-training models primarily focus on predicting a diverse range of physical quantities, including torsion angles, angles between edges, alpha carbon positions, and distances, as well as residue types. (Zhang et al., 2022b; Guo et al., 2022; Chen et al., 2023b; Wang et al., 2022b). Contrastive learning has demonstrated exceptional performance in a recent study based on GearNet (Zhang et al., 2022b). Additionally, recent attempts for pre-training with protein surfaces have been made (Lee et al., 2023). Numerous diffusion-based models have been proposed (Huang et al., 2023; Zhang et al., 2023). For instance, SiamDiff (Zhang et al., 2023) employs a pre-trained protein encoder through sequence-structure joint diffusion modeling. HotProtein(Chen et al., 2022), a structure-aware pre-training model, is proposed to improve thermostability prediction. A self-supervised pretraining method is also used to predict compound–protein affinity and contact(You & Shen, 2022).
@@ 3.1 Vector Aware Bilevel Sparse Attention Network
Protein Graph Construction We first construct both residue-level and atom-level graphs. A residue level graph can be defined as $\mathcal{G}_{res}=(\mathcal{V}_{CA},\mathcal{E}_{res})$, where $\mathcal{V}_{CA}$ includes all alpha carbon nodes and $\mathcal{E}_{res}$ includes all residue edges. An atom level graph can be defined as $\mathcal{G}_{atom}=(\mathcal{V},\mathcal{E}_{atom})$, where $\mathcal{V}$ includes all-atom nodes, i.e. $\mathcal{V}_{CA}\subset\mathcal{V}$ and $\mathcal{E}_{atom}$ includes all atom level edges. The construction of edges involves the selection of k nearest neighbors in the space for both residue and atom-level graph nodes. The residue-level graph shares alpha carbon nodes with the atom-level graph. Consequently, the bilevel graph of protein can be written as $\mathcal{G}=(\mathcal{V},\mathcal{E}_{res},\mathcal{E}_{atom})$. The model architecture can be seen in Figure2. This approach facilitates information exchange between atom and residue levels, thereby not only allowing atoms to acquire an expanded receptive field but also enabling residues to obtain detailed structural information.
To comprehensively encapsulate global information on protein, we introduce a virtual origin point positioned at the geometry center. This point is connected with every atom and residue, integrating overall protein representation.
Node Encoding In our model, node embeddings incorporate both atom type and residue type. Additionally, we utilize the Large language model (LLM) version ESM (Lin et al., 2023) to leverage sequence information. The representation of atom nodes can be seen as follows:
@@
where $ATOM$ and $RES$ stand for the type of atom and residue. $\boldsymbol{W_{E}}$ maps the shape of ESM repr. to node shape.
Edge Distance Encoding To encode the distance between atoms or residues, we use the Gaussian kernel.
@@
Edge Vector Encoding In previous protein-pretraining modeling, structural information is traditionally encoded using the distance between residues. While this distance-based structural encoding may suffice for residue modeling, encoding distance is not informative enough for atom modeling since the number of atoms is much larger than that of the residueB. Encoding edge directions in both the residue local coordinate system and the absolute global coordinate system can alleviate this problem. The residue local coordinate system can be formulated as following steps with three backbone atoms(CA, C, N) in each residue. We use $\boldsymbol{r}$ to indicate the position of atoms and set $\boldsymbol{v}_{N}=\boldsymbol{r}_{N}-\boldsymbol{r}_{CA}$ and $\boldsymbol{v}_{C}=\boldsymbol{r}_{C}-\boldsymbol{r}_{CA}$ as the vector of edges between N and CA, CA and C in the absolute global coordinate system. More generally, We define $\boldsymbol{v}_{l}$ and $\boldsymbol{v}_{g}=\boldsymbol{r}_{atom}-\boldsymbol{r}_{CA}$ are the vectors of edges from CA to an atom in the same residue in the local residue coordinate system and global coordinate system, respectively. They serve to encode edge directions in subsequent steps. Then we can get a local coordinate system with a rotational matrix $\boldsymbol{R}$, which is used to transform vectors in global coordinates to local coordinates. $\boldsymbol{R}$ is constructed with unit vector $\boldsymbol{u},\boldsymbol{v},\boldsymbol{w}$ in XYZ axis of residue local coordinate.
@@
where $\boldsymbol{l}(i-j)$ is the learnable positional encoding based on the sequence distance between two residues. $\boldsymbol{W_{g}}$ and $\boldsymbol{W_{f}}$ maps the dimension of distance and direction encoding to dimension of edge.
Sparse Attention Module(SAM) Numerous pre-training models for molecules based on Graphormer have achieved state-of-the-art performance in downstream tasks (Zhou et al., 2023). However, a fully connected Graphormer model can pose significant computational challenges (Lu et al., 2023). To tailor Graphormer’s attention mechanism to the all-atom level graph, we introduce a Sparse Graph Attention Module aimed at reducing computational load and preventing overfitting in downstream tasks. The sparse attention module is formulated as follows:
@@
where $\boldsymbol{x}^{l}_{i}$ is the representation of node $i$ in the $l$th layer. $a^{l}_{i,j}$ is the attention weight. Set $\mathcal{N}_{i}$ includes all nodes connected to node $i$. $\boldsymbol{W_{B}}$ is a linear layer for bias calculation. $\boldsymbol{e}_{i,j}$ is the edge representation between node $i$ and node $j$. The $h$ dimension in multi-head attention is omitted.
@@ 3.2 Pre-training Tasks
Span Mask Protein Chain(SMPC) Residue Type Prediction During the pre-training phase, span masking is applied to mask residue types of consecutive biologically meaningful substructures. Specifically, only the alpha carbon atom is retained for the span-masked residues, while all other atoms are eliminated. This aims to prevent information leakage from side chain atoms and other backbone atoms, which could make residue type prediction and torsion angle prediction trivial. We find even backbone atoms can leak residue type to some extent. Subsequently, only alpha carbon representation is employed for residue type prediction. Cross-entropy loss serves as the objective function for predicting residue types.
Span Mask Protein Chain(SMPC) Side Chain and Backbone Torsion Angle Prediction Drawing upon the feature of alpha carbon atoms in SMPC part, Vabs-Net predicts the cosine and sine components for seven side chain and backbone torsion angle angles. In addition to the L1 norm loss for cosine and sine values, the loss function encompasses a term that ensures the normalization of the sum of squared sine and cosine values to promote accurate angular representation. The loss function for this task is similar to the one used in AlphaFold2 (Jumper et al., 2021).
Atom Position and Distance Prediction Apart from the residue level pre-training tasks. To learn atom-level structural information, We randomly choose consecutively residues independent of SMPS. For these residues, we simply add Gaussian noise to atoms in those residues instead of removing atoms. For these atoms, positions are predicted by a movement prediction head.
@@
Solvent Accessible Surface Area(SASA) Prediction SASA describes the surface area of a biomolecule, such as a protein or nucleic acid, that is accessible to a solvent. It could be used to understand the shape of a protein. We used freesasa (Mitternacht, 2016) to calculate SASA of each atom. The loss function for SASA prediction is L1 loss.
@@ 4 Experiments
In this section, we report our experiment setup and results in training and evaluation of our models for pre-training and downstream tasks. Detailed information about downstream task settings can be seen in Appendix D.
@@
Method BP MF CC EC
MTL(Wang et al., 2022b) 0.445 0.640 0.503 0.869
GRADNORM(Chen et al., 2018; Wang et al., 2022b) 0.466 0.643 0.504 0.874
LM-GVP(Wang et al., 2022a) 0.417 0.545 0.527 0.664
ROTOGRAD(Javaloy & Valera, 2022) 0.470 0.638 0.509 0.876
CDConv(Fan et al., 2022) 0.453 0.654 0.479 0.820
PROMPTPRO.(Wang et al., 2022b) 0.495 0.677 0.551 0.888
ESM_1b(Lin et al., 2023) 0.470 0.657 0.488 0.864
GearNet(Zhang et al., 2022b) 0.490 0.650 0.486 0.874
Siamdiff(Zhang et al., 2023) - - - 0.857
GearNet-ESM(Zhang et al., 2022b) 0.516 0.684 0.506 0.890
SiamDiff-ESM(Zhang et al., 2023) - - - 0.897
GearNet-ESM-INR-MC(Lee et al., 2023) 0.518 0.683 0.504 0.896
Vabs-Net-no-SMPC 0.496 0.667 0.552 0.876
Vabs-Net 0.531 0.695 0.579 0.900
@@ 4.3 Protein Binding Site Prediction
Method pre-train B277 DT198 ASTEX85 CHEN251 COACH420
FPocket(Le Guilloux et al., 2009) $\times$ 31.5 23.2 34.1 25.4 30.0
SiteHound(Hernandez et al., 2009) $\times$ 36.4 23.1 38.9 29.4 34.9
MetaPocket2(Macari et al., 2019) $\times$ 37.3 25.8 37.5 32.8 37.7
DeepSite(Jiménez et al., 2017) $\times$ 34.0 29.1 37.4 27.4 33.9
P2Rank(Krivák & Hoksza, 2018) $\times$ 49.8 38.6 47.4 56.5 45.3
ESM2_150M(Lin et al., 2023) $\surd$ 19.6 16.6 20.5 18.9 22.0
GearNet(Zhang et al., 2022b) $\surd$ 39.9 35.8 41.0 36.4 41.3
Siamdiff(Zhang et al., 2023) $\surd$ 37.7 31.0 40.7 35.3 40.3
Vabs-Net $\times$ 57.7 48.6 57.8 53.2 61.4
Vabs-Net $\surd$ 60.1 52.0 58.8 56.3 64.1
@@
Method pre-train auc
TargetDNA(Hu et al., 2016) $\times$ 0.825
DNAPred(Zhu et al., 2019) $\times$ 0.845
SVMnuc(Su et al., 2019) $\times$ 0.812
COACH-D(Wu et al., 2018) $\times$ 0.761
NucBind(Su et al., 2019) $\times$ 0.797
DNABind(Liu & Hu, 2013) $\times$ 0.858
@@
ESM2_150M(Lin et al., 2023) $\surd$ 0.779
GearNet(Zhang et al., 2022b) $\surd$ 0.849
Siamdiff(Zhang et al., 2023) $\surd$ 0.823
Vabs-Net $\times$ 0.912
@@
Setting We have constructed a large high-quality small molecule binding site dataset, and we use this dataset to train our model and other baselines, which will be open access to the public. For all-atom models, the label of Alpha C is the residue label. In DNA and RNA binding site prediction, we use the AUC for evaluation. Baselines are finetuned using the same setting as their original paper.
In the preparation of our small molecule training dataset, we utilize three distinct datasets: CASF-2016 coreset, PDBBind v2020 refined set, and MOAD. During dataset preparation, we follow the original ligand records in these datasets to extract the protein and ligand components based on the corresponding PDB ID from RCSB. The extracted segments undergo a structural repair process using in-house scripts. For proteins, this process includes the repair of missing residues, replenishment of absent heavy atoms, and addition of hydrogen atoms. In the case of ligands, the process involves repairing bond orders, adding hydrogen atoms, and determining the correct protonation states based on the pocket environment. A comparison of our dataset and frequently used scPDB can be seen in Appendix A. To avoid leakage, MMSeqs2 is used to filter out high protein sequence similarity(similarity above 40% which is also used in AlphaFold2 (Jumper et al., 2021)) with test sets.